OpenStudy (anonymous):

Is my answer correct? Solve. 3x2 - 2x - 3 = 0 x= 1 _+_ sq root 10 over 3

OpenStudy (anonymous):

Answer looks correct.

OpenStudy (youngster):

yeah, it seems correct.

OpenStudy (anonymous):

. 3x^2 – 2x – 3 = 0 USe the quadratic equation where a = 3, b = -2, and c = -3 x = (-b +/- sqrt(b^2 - 4ac))/(2a) x = (2 +/- sqrt((-2)^2 - 4(3)(-3))/(2(3)) x = (2 +/- sqrt(4 + 36))/6 x = (2 +/- sqrt(40))/6 x = (2 +/- sqrt(4 * 10))/6 x = (2 +/- 2* sqrt(10))/6 x = (1 +/- sqrt(10))/3 ANSWER so your answer is correct

OpenStudy (anonymous):

thank you guys! means so much!